OpenStudy (anonymous):

Melissa returned home from mowing lawns at 4:15 P.M. It took 1 3/4 h to mow the first lawn. The second and third lawns she mowed each took 1 1/5 h. She took only a 45-min break. When did Melissa begin mowing the first lawn?

OpenStudy (anonymous):

12:33 P.M

OpenStudy (anonymous):

no

OpenStudy (anonymous):

11:21 A.M

OpenStudy (mathstudent55):

1.75 hour for the first lawn 1.2 hour for second lawn 1.2 hour for third lawn 0.75 hour break Total time of mowing lawns + break = 4.9 hours = 4 hours and 54 minutes What is 5 hours before 4:15 pm? Then add 6 minutes to that.

OpenStudy (anonymous):

1 h = 60 min 1 3/4 = 60 + 45 = 95 min 1 1/5 = 60 + 12 = 72 min 95 min + 72 min + 72 min = 293 min = 4,88h 4h = 4h 0,88 h = 52 min 4.15PM - 4h. 52 min = is the solution